OpenStudy (anonymous):

a cylinder has a radius of 8 inches and a height of 26 inches. what is the approximate LATERAL area?

OpenStudy (rsadhvika):

Lateral Area = \(2 \pi r h \) \(r\) = radius \(h\) = height

OpenStudy (jhannybean):

The formula forlateral area of a cylinder is \(\large LA=2\pi rh\)\[\large r= 8 \ , \ h= 26\]\[\large LA= 2(\pi)(8)(26) \ in^2 \]

OpenStudy (anonymous):

rsadhvika u have been a boss u know everything......how do you give people medals

OpenStudy (jhannybean):

Thank you.

OpenStudy (anonymous):

yep

OpenStudy (rsadhvika):

np :) i was a born boss lol :P